# Pattern formation in a coupled driven diffusive system

### v1

@@ IV Numerical integration of the Stochastic Field Theory
In this section, we integrate Eq. (III) numerically to study how the possible steady states depend on the various parameters of the model. Our goal in this section is to assess whether the continuum stochastic field theory we propose in Eq. (III) is capable of reproducing the different regimes encountered both in the DWRLG and the FLM, namely the microemulsion structure in the low-density phase, and stripes aligned perpendicular to drive in the high-density phase. Additionally, we check whether the continuum model also exhibits the irregular stripe regime present in the FLM.
To numerically integrate Eq. (III) in 2D with periodic boundary conditions, we design a numerical SPDE solver using the PSSETD2H and smooth dealiasing. To fix the timestep $\Delta t$, we run simulations at several values of $\Delta t$ and choose the largest value such that there are no appreciable changes in the results; in practice we use $\Delta t=0.5$. Unless otherwise specified, simulations are performed on a domain of size $L=128$, using grid resolutions of $N=200$ points per direction. We also adopt the following initial conditions for the fields $\phi_{\pm}$:
@@
where $\sigma_{0}=10^{-2}$ and $\xi(\mathbf{r})$ is a conserved Gaussian noise with zero mean and unit correlation.
@@
In the low-density equilibrium regime, $\rho<\rho_{c}^{\mathrm{FT}}=0.37\dots$ and $\delta=0$, for $\lambda\geq 0$, both the CF and DF steady-state structure factors are isotropic and follow the usual Ornstein-Zernike form. For a fixed drive $\delta$ in this regime, exploring the full phenomenology of the model requires scanning a four-dimensional parameter space (or five, if the noise amplitudes are included), which is computationally demanding. Fortunately, we can simplify the task since our main goal is to find the various phases of the lattice models. Thus, first, at low densities $\rho$, the non-linearities facilitating coarsening dynamics should be irrelevant (as $D_{-}>0$), so we set $\lambda=0$. Then, taking advantage of the perturbative analysis in Ref. [2], we set $g_{0}=4.243$, $g_{-}=0.707$, $g_{+}=2.828$, and $\sigma=0.27$, which should correspond to the microemulsion phase. Fig. 8 shows the steady-state charge fluctuation SF, $S_{-}(\mathbf{q})$, in the resulting microemulsion phase, averaged over $5000$ time points evenly spaced across the final $10^{6}$ time steps, and using $150$ independent realizations. We observe, as expected, both a discontinuity at the origin and the emergence of a peak at a nonzero wavevector, $q^{*}=0.190\pm 0.009$, along the drive direction, closely matching the behavior of the FLM. The peak is well fit by the same fitting form used for the FLM, $A_{1}\,q^{2}\left[A_{2}\,q^{2}+q^{2}(q-q^{*})^{2}\right]^{-1}$. These results are consistent with previous studies [2], where it is also noted that the field theory smoothes the peak region, making it noticeably less sharp than in the lattice-based models.
@@
Figure 8: Microemulsion phase for the stochastic equations of motion; Eq. (III). Linear coefficients are fixed by Eq. (III), using $\rho=0.2,$ and $\delta=1$. The noise intensities are fixed to $\sigma=0.27$. The other parameters are retrieved from [2] using the same $\rho$ and $\delta$: $\lambda=0,\,\,g_{0}=4.243,\,g_{-}=0.707$ and $g_{+}=2.828$. (a) Charge fluctuation (CF, black circles) and density fluctuation (DF, orange diamonds) structure factors (SFs) along the drive direction. The error bars are smaller than the symbols. Blue line is the best fit $A_{1}\,q^{2}\left[A_{2}\,q^{2}+q^{2}(q-q^{*})^{2}\right]^{-1}$, using $10$ points around the peak. We find $q^{*}=0.190\pm 0.009$. Note the shoulder in the DF SF at $2q^{*}$. (b) CF SF. Since the peak near the origin is difficult to discern from just the color map, we also mark the contour of the levels at $0.8,0.9,0.95,0.97,0.99$ and $0.999\cdot S_{-}(q^{*},0)\ (\simeq 0.253)$ by black dashed lines to guide the eyes. (c) CF and (d) DF typical steady-state field configurations. This figure is analogous to Fig. 6 for the FLM (and Fig. 10 in [2] for the DWRLG).
@@
Another simple case is the high-density equilibrium regime, $\rho>\rho_{c}^{\mathrm{FT}}$ and $\delta=0$. Here, all the drive-dependent couplings vanish and we simply set a positive value for $\lambda$. In this case, a steady-state amplitude is systematically found  [^12] This is supported by the weakly nonlinear stability analysis of Eq. (III) [48], which can also be used to crudely estimate the amplitude values slightly above the onset, $\rho\gtrsim\rho_{c}^{\mathrm{FT}}$. On the other hand, for $\lambda<0$, Eq. (III) has no bounded solutions and the numerical integration is ill-behaved, thus making necessary the inclusion of higher-order derivative terms. for the charge and density fluctuations. In all the runs performed, we consistently find two types of steady-state configurations: a round domain that resembles a bubble (see Fig. 9), or a single stripe whose orientation varies with realization. In either case, there is a single domain dominated by each species, which is consistent with the results of the FLM in this undriven regime. Furthermore, we verify that our equations of motion follow the LS law at high densities ($\rho>\rho_{c}^{\mathrm{FT}}$), with the growth of the SF peaks scaling as $t^{1/3}$, so that the characteristics of the coarsening dynamics do not depend on the particular value of $\lambda$. In fact, the cubic saturation terms due to $\lambda$ are more relevant to fix the magnitude of the steady-state amplitude and, as a result, their actual values may depend on the choice of $\lambda$. In the context of the CH equation or Model B [59, 60, 61], these couplings are often set to unity, which corresponds to a rescaling of the field values. Therefore, for simplicity, we fix $\lambda=1.0$ for the remainder of this section.
@@
Our largest parameter space corresponds to the driven, high-density regime. This time, we cannot rely on the coupling values contained in Ref. [2], as they are specialized for the low-density phase. However, we can reduce the parameter space by discarding the term $g_{+}\phi_{+}^{2}$ in Eq. (III); see the simplifying assumptions in Sec. III. We then proceed by fixing a reasonable value for the noise strengths, $\sigma$. Fig. 9 compares typical runs with and without the external conserved noises $\xi_{\pm}$ in (III) for $\rho=0.7$, $\delta=0$; and $\rho=0.7$, $\delta=1.0$, $g_{-}=0.5$, $g_{0}=0$. The linear coefficients are given by the expressions in (III). In accordance with observations made for models A and B in [59], we note that the external additive noise has little effect on the behavior  [^13] This is not true, however, for multiplicative noises [59].. We do expect the noise to become more relevant in the low-density regime and at interfaces, as we have verified for the microemulsion phase  [^14] In fact, for $\rho<\rho_{c}^{\mathrm{FT}}$, the fields will decay to zero, even if $\delta$ is maximal. Therefore, the noises act as source terms, which are crucial to the development of the microemulsion behavior. In another words, the microemulsion phase cannot be observed without the noise terms in Eqs. (III).. Since no appreciable quantitative changes are found when varying $\sigma$ (within some range) in the ordered phase, we fix the noise intensities to a moderate value $\sigma=10^{-2}$.
@@
Figure 9: Comparison of the charge fluctuation steady-state field configuration with and without the additive noise in Eq. (III) for $L=128$, $\rho=0.6$ and $\lambda=1.0$. From left to right, the effective noise intensities (Eq. (16)) are $\sigma=0$ (a), $0.01$ (b) and $0.03$ (c) $0.04$ (d). Upper plots: $\delta=0$. Lower plots: $\delta=1,g_{0}=0,g_{-}=0.5$. No appreciable quantitative differences in the results are found for $\sigma=0,0.1$, in agreement with results for Model A and B in [59]. Raising the noise strength further makes the dynamics noise-dominated.
@@
Therefore, with all these simplifications, we are left with two nonlinear couplings to consider: $_{0}, gg_{-}$. To sweep the parameter space, we shall limit ourselves to $g_{0},g_{-}\in[0,1]$, which is a typical range indicated by the different coarse-graining prescriptions in the high-density region. We also vary $\Delta v$, which depends on the drive and density as shown in Eq. (III). Figure 10 summarizes the four different regimes we encounter by performing a series of numerical integrations:
@@
    We find that for $0<\Delta v<\Delta v^{*}$, vertical stripes are not formed in all simulations performed. This supports the conclusion that while $\Delta v>0$ is a necessary condition for stripe formation with the correct alignment, it is not sufficient: $\Delta v$ must exceed a critical threshold $\Delta v^{*}$ to stabilize the vertical stripe pattern.
@@
Figure 10: Typical configurations of the field equations for the regimes (1)-(4) mentioned in the text. Upper(lower) plots are the charge fluctuation (density fluctuation) field configurations. We set $\rho=0.7$, fixing all drive-independent linear coefficients, $D_{\pm}$, $\Gamma_{\pm}$ in Eqs. (III) via the GE; see Eq. (III). In (1)-(3) we use $\delta=1$ and (1) $g_{-}=g_{0}=0$; (2) $g_{-}=0.5$ and $g_{0}=0$; (3) $g_{-}=0.5$ and $g_{0}=0.05>g_{0}^{*}$. In (4), $\Delta v=0.05<\Delta v^{*}$ (corresponding to $\delta=0.24$), while $g_{-}=0.5$ and $g_{0}=0.0$.
@@
Our numerical results suggest that the term proportional to $g_{-}$ in Eq. (III) is necessary to stabilize the ordered stripes, analogous to the drive-independent stability term proportional to $\lambda$. Thus, like $\lambda$, we will fix the value of $g_{-}$ in the following. The choice to fix $g_{-}$ is further supported by the observation that, for finite $\Delta v>\Delta v^{*}$, we verified that keeping $g_{-}\lesssim 0.5$ does not affect the stripe orientation. Additionally, our simulations show that increasing $g_{0}$ consistently destabilizes perpendicular stripes, moving the system away from the phenomenology of the lattice models. Accordingly, we set $g_{-}=0.5$ and $g_{0}=0$ in what follows  [^15] Note that it is not strictly necessary to drastically set $g_{0}=0$; any small value $g_{0}\approx 0$ would yield similar results. However, since no particular value is preferred, we choose zero for simplicity and computational efficiency.. In other words, we focus the remainder of our studies on Regime 2 shown in Fig. 10. In this regime, we explore the effect of varying $\rho$ [which will vary the linear terms shown in Eq. (III)] and $\Delta v$ (or equivalently $\delta$), and find that increasing either parameter also increases order along the drive direction. Conversely, when $\rho$ and $\Delta v$ are close to the thresholds $(\rho_{c}^{\mathrm{FT}},\Delta v^{*})$, we find evidence of long-wavelength (Eckhaus-like) instabilities within the perpendicular stripe regime, which act as a mechanism to reduce the number of stripes, similarly to what occurs in the FLM. Based on this, we identify these two as the regular and irregular stripe regimes, respectively. The resemblance goes beyond visual inspection of the configurations: we also detect the clear signatures of the different regimes in the charge SF, with the irregular stripe regime [Fig. 11(a)] showing a broad peak near the origin, whereas the regular stripe regime [Fig. 11(b)] displays a sharp, Dirac delta-like peak accompanied by pronounced higher-order harmonics. However, unlike the FLM irregular stripe regime, we observe weak harmonics around the peak and secondary maxima at $q_{x}>q^{*}$ for $q_{y}\neq 0$, indicating stripes that are more ordered along $x$ yet weakly modulated along  [^16] In fact, this vertical modulation seems to be a feature of the equations that is already present if $g_{\pm,0}=0$; see, e.g., Fig. 10  (1). $y$. Additionally, in the regular stripe regime, the continuum model further accentuates ordering: the SF is heavily localized at $q_{y}=0$, with modes at any $q_{y}\neq 0$ appearing almost as background noise; the resulting spectrum is strongly peaked and displays deep dips, resembling Bragg peaks in a diffraction grating that are consistent with a highly ordered stripe pattern.
@@
For example, comparing with the results in Table 1, the ratio of peak values $S_{-}(q^{*},q_{y})$ for $q_{y}L/2\pi=0,1$ in the continuum model using $\delta=1$ is $S_{-}\left(q^{\ast},0\right)/S_{-}\left(q^{\ast},0.05\right)\simeq 52$ in the irregular stripe regime ($\rho=0.4$) and $S_{-}\left(q^{\ast},0\right)/S_{-}\left(q^{\ast},0.05\right)\simeq 2700$ in the regular stripe regime ($\rho=0.6$) (nearly an order of magnitude larger than the FLM!). Table 2 lists the peak values for the first three $q_{y}$ modes available in these two cases. In addition, we note that the density at which the regular stripe regime emerges in the continuum model appears to be shifted relative to the FLM: regular stripes are observed already at $\rho=0.6$ for $\delta=1$ (whereas in the FLM, $\rho_{u}\simeq 0.65$). However, note that our $\rho$-dependence in the continuum model only refers to the GE results for the linear coefficients [see Eq. (III)]. These assumptions are violated at high densities and we generally expect the location of the transition to be shifted due to the nonlinearities and noise.
@@
Figure 11: Phases of irregular stripes and regular stripes found by numerical integration of Eq. (III), analogous to the ordered phases of the FLM illustrated in Fig. 5. We use $\delta=1$, maintaining $\lambda=1,\,g_{0}=0$ and $g_{-}=0.5$. (a) $\rho=0.4$. (b) $\rho=0.6$, with linear coefficients fixed by Eqs. (III). Charge fluctuation (left panels) and density fluctuation (center panels) typical steady-state field configurations. Right panels: charge fluctuation structure factor, $S_{-}(q_{x},q_{y})$, as a function of $q_{x}$ for $q_{y}\,L/2\pi=0,...,10$, in the two regimes.
@@
Finally, by fixing $\rho$ and varying the drive-dependent parameters $\Delta v$, $g_{0}$, and $g_{-}$, we uncover a wide range of regimes in the SFT, many of which have not been reported in the previous lattice-based studies. These regimes largely occur when $g_{0}\phi_{+}\gg\Delta v$. To (partially) understand the origin of these new regimes, we reorganize the drive-dependent terms proportional to $\partial_{x}\phi_{-}$ in Eq. (III) into the combination
@@
If $g_{0}\phi_{+}\gg\Delta v$, we see from Eq. (21) that the flow direction of the CF is strongly dependent on the sign of $\phi_{+}$, facilitating the formation of vortices and complex flow patterns.
Fig. 13 shows snapshots taken at long times $t=10^{6}$ that illustrates the variety of new behaviors: regular [Fig. 13(a)] and irregular [Fig. 13(b)] stripes aligned with the drive, spatio-temporal chaos [Fig. 13(c)] and phase separation into bubbles modulated by shock-like envelopes [Fig. 13(d)]. Because these regimes do not appear in either the FLM or the DWRLG, we have restricted our primary analysis to parameter regions that better reflect those models. Nevertheless, Fig. 13 clearly demonstrates that even a minimal set of continuum equations can generate a remarkably rich phenomenology. While a systematic exploration of how each term contributes to the various steady states would be valuable, such an endeavor is computationally demanding and falls beyond the scope of this study. It is worth noting that similar regimes have been observed in other contexts, such as parallel stripes in the KLS model [67] and spatio-temporal chaos in the active non-reciprocal Cahn-Hilliard model [30].
Figure 13: Charge fluctuation $\phi_{-}$ snapshots at $t=10^{6}$ of the different regimes found by numerical integration of Eq. (III). We use $\rho=0.7$ [fixing the coefficients in Eq. (III)], $\lambda=1$, and $\sigma=0.01$ in all panels. (a) Regular and (b) irregular stripes aligned with the drive. (c) Spatio-temporal chaos. (d) Phase separation into bubbles modulated by shock-like envelopes. Parameters: (a) $\Delta v=0.05$, $g_{0}=0.003$, $g_{-}=0.5$. (b) $\Delta v=0.211$, $g_{0}=0.3$, $g_{-}=0.5$. (c) $\Delta v=0.211$, $g_{0}=0.03$, $g_{-}=0.5$. (d) $\Delta v=0.05$, $g_{0}=0.1$, $g_{-}=0.5$.

### v2

@@ IV Numerical integration of the Stochastic Field Theory
In this section, we integrate Eq. (III) numerically to study how the possible steady states depend on the various parameters of the model. Our goal in this section is to assess whether the continuum stochastic field theory we propose in Eq. (III) is capable of reproducing the different regimes encountered both in the DWRLG and the FLM, namely the microemulsion structure in the low-density phase, and stripes aligned perpendicular to the drive in the high-density phase. Additionally, we check whether the continuum model also exhibits the irregular stripe regime present in the FLM.
To integrate Eq. (III) numerically in 2D with periodic boundary conditions, we design a numerical SPDE solver using the PSSETD2H and smooth dealiasing. To fix the timestep $\Delta t$, we run simulations at several values of $\Delta t$ and choose the largest value such that there are no appreciable changes in the results; in practice we use $\Delta t=0.5$. Unless otherwise specified, simulations are performed on a domain of size $L=128$, using grid resolutions of $N=200$ points per direction  [^10] This mesh size is chosen to balance numerical accuracy and computational cost. We have verified that further refining the mesh does not produce any appreciable changes in the results.. We also adopt the following initial conditions for the fields $\phi_{\pm}$:
@@
where $\sigma_{0}=10^{-2}$ and $\xi(\mathbf{r})$ is a conserved Gaussian noise with zero mean and unit variance.
@@
In the low-density equilibrium regime, $\rho<\rho_{c}^{\mathrm{FT}}=0.37\dots$ and $\delta=0$, for $\lambda\geq 0$, both the CF and DF steady-state structure factors are isotropic and follow the usual Ornstein-Zernike form. For a fixed drive $\delta$ in this regime, exploring the full phenomenology of the model requires scanning a four-dimensional parameter space (or five, if the noise amplitudes are included), which is computationally demanding. Fortunately, we can simplify the task since our main goal is to find the various regimes of the lattice models. Thus, first, at low densities $\rho$, the non-linearities facilitating coarsening dynamics should be irrelevant (as $D_{-}>0$), so we set $\lambda=0$. Then, taking advantage of the perturbative analysis in Ref. [2], we set $g_{0}=4.243$, $g_{-}=0.707$, $g_{+}=2.828$, and $\sigma=0.27$, which should correspond to the microemulsion regime. Fig. 8 shows the steady-state charge fluctuation SF, $S_{-}(\mathbf{q})$, in the resulting microemulsion phase, averaged over $5000$ time points evenly spaced across the final $10^{6}$ time steps, and using $150$ independent realizations. As expected, $S_{-}(\mathbf{q})$ exhibits both a discontinuity at the origin and the emergence of a peak at a nonzero wavevector, $q^{*}=0.190\pm 0.009$, along the drive, closely matching the behavior of the FLM. The peak is well fit by the same fitting form used for the FLM, $A_{1}\,q^{2}\left[A_{2}\,q^{2}+q^{2}(q-q^{*})^{2}\right]^{-1}$. These results are consistent with previous studies [2], where it was also noted that the field theory smoothes the peak region, making it noticeably less sharp than in the lattice-based models.
@@
Figure 8: Microemulsion phase for the stochastic equations of motion; Eq. (III). Linear coefficients are fixed by Eq. (B), using $\rho=0.2,$ and $\delta=1$. The noise intensities are fixed to $\sigma=0.27$. The other parameters are retrieved from [2] using the same $\rho$ and $\delta$: $\lambda=0,\,\,g_{0}=4.243,\,g_{-}=0.707$ and $g_{+}=2.828$. (a) Charge fluctuation (CF, black circles) and density fluctuation (DF, orange diamonds) structure factors (SFs) along the drive direction. The error bars are smaller than the symbols. Blue line is the best fit $A_{1}\,q^{2}\left[A_{2}\,q^{2}+q^{2}(q-q^{*})^{2}\right]^{-1}$, using $10$ points around the peak. We find $q^{*}=0.190\pm 0.009$. Note the shoulder in the DF SF at $2q^{*}$. (b) CF SF. Since the peak near the origin is difficult to discern from just the color map, we also mark the contour of the levels at $0.8,0.9,0.95,0.97,0.99$ and $0.999\cdot S_{-}(q^{*},0)\ (\simeq 0.253)$ by black dashed lines to guide the eyes. (c) CF and (d) DF typical steady-state field configurations. This figure is analogous to Fig. 6 for the FLM (and Fig. 10 in [2] for the DWRLG). A movie related to this figure is available in [46].
@@
Another simple case is the high-density equilibrium regime, $\rho>\rho_{c}^{\mathrm{FT}}$ and $\delta=0$. Here, all the drive-dependent couplings vanish and we simply set a positive value for $\lambda$. In this case, a steady-state amplitude is systematically found  [^11] This is supported by the weakly nonlinear stability analysis of Eq. (III) [43], which can also be used to crudely estimate the amplitude values slightly above the onset, $\rho\gtrsim\rho_{c}^{\mathrm{FT}}$. On the other hand, for $\lambda<0$, Eq. (III) has no bounded solutions and the numerical integration is ill-behaved, thus making necessary the inclusion of higher-order derivative terms. for the charge and density fluctuations. In all the runs performed, we consistently find two types of steady-state configurations: a round domain that resembles a bubble (see Fig. 9), or a single stripe whose orientation varies with realization. In either case, there is a single domain dominated by each species, which is consistent with the results of the FLM in the undriven regime. Furthermore, we verify that our equations of motion follow the Lifshitz-Slyozov law at high densities ($\rho>\rho_{c}^{\mathrm{FT}}$), with the growth of the SF peaks scaling as $t^{1/3}$, so that the characteristics of the coarsening dynamics do not depend on the particular value of $\lambda$. In fact, the cubic saturation terms due to $\lambda$ are more relevant to fix the magnitude of the steady-state amplitude and, as a result, their actual values may depend on the choice of $\lambda$. In the context of the Cahn-Hilliard equation or Model B [54, 55, 56], these couplings are often set to unity, which corresponds to a rescaling of the field values. Therefore, for simplicity, we fix $\lambda=1.0$ for the remainder of this section.
@@
The driven, high-density regime involves a large parameter space. Here, we cannot rely on the coupling values obtained in Ref. [2], as they are specialized for the low-density phase. However, we can reduce the parameter space by discarding the term $g_{+}\phi_{+}^{2}$ in Eq. (III); see the simplifying assumptions in Appendix B. We then proceed by fixing a reasonable value for the noise strengths, $\sigma$. Fig. 9 compares typical runs with and without the conserved noises $\xi_{\pm}$ in (III) for $\rho=0.7$, $\delta=0$; and $\rho=0.7$, $\delta=1.0$, $g_{-}=0.5$, $g_{0}=0$. The linear coefficients are given by the expressions in (B). We find that quantitative properties in the ordered phase are largely insensitive to the noise amplitude within a moderate range  [^12] This is not true, however, in the low-density regime and at interfaces, as we have verified for the microemulsion phase. In fact, for $\rho<\rho_{c}^{\mathrm{FT}}$, the fields will decay to zero, even if $\delta$ is maximal. Therefore, the noises act as source terms, which are crucial to the development of the microemulsion behavior. In another words, the microemulsion phase cannot be observed without the noise terms in Eqs. (III)., consistent with observations for Models A and B in [54]. Accordingly, we fix the noise intensity to a moderate value, $\sigma=0.01$. At very high noise levels, however, the system displays qualitatively new behaviors, most notably stripe reorientation [see Fig. 9(c,d)], suggesting that the high-noise limit of the model may contain additional phenomena worthy of future investigation.
@@
Figure 9: Comparison of the charge fluctuation steady-state field configuration with and without the additive noise in Eq. (III) for $L=128$, $\rho=0.6$ and $\lambda=1.0$. From left to right, the effective noise intensities (Eq. (15)) are $\sigma=0$ (a), $0.01$ (b), $0.03$ (c) and $0.04$ (d). Upper plots: $\delta=0$. Lower plots: $\delta=1,g_{0}=0,g_{-}=0.5$. No appreciable quantitative differences in the results are found for (a) and (b), in agreement with results for Model A and B in [54]. Raising the noise strength further makes the dynamics noise-dominated and promotes stripe reorientation.
@@
Therefore, following these simplifications, we are left with two nonlinear couplings to consider: $g_{0}$ and $g_{-}$. To sweep the parameter space, we shall limit ourselves to $g_{0},g_{-}\in[0,1]$, which is a typical range indicated by the different coarse-graining prescriptions in the high-density regime. We also vary $\Delta v$, which depends on the drive and density as shown in Eq. (B). Figure 10 summarizes the four different regimes we encounter via numerical integration:
@@
    For $0<\Delta v<\Delta v^{*}$, vertical stripes are not formed in any simulation. This supports the conclusion that while $\Delta v>0$ is a necessary condition for stripe formation with the correct alignment, it is not sufficient: $\Delta v$ must exceed a critical threshold $\Delta v^{*}$ to stabilize the vertical stripe pattern.
@@
Figure 10: Typical configurations of the field equations for regimes (1)-(4) mentioned in the text. Drive to the right. Upper (lower) plots are the charge fluctuation (density fluctuation) field configurations. We set $\rho=0.7$, fixing all drive-independent linear coefficients, $D_{\pm}$, $\Gamma_{\pm}$ in Eqs. (III) via the GE; see Eq. (B). In (1)-(3) we use $\delta=1$ and (1) $g_{-}=g_{0}=0$; (2) $g_{-}=0.5$ and $g_{0}=0$; (3) $g_{-}=0.5$ and $g_{0}=0.05>g_{0}^{*}$. In (4), $\Delta v=0.05<\Delta v^{*}$ (corresponding to $\delta=0.24$), while $g_{-}=0.5$ and $g_{0}=0.0$. A movie related to this figure is available in [46].
@@
Our numerical results suggest that the term proportional to $g_{-}$ in Eq. (III) is necessary to stabilize ordered stripes, analogous to the drive-independent stability term proportional to $\lambda$. Thus, like $\lambda$, we will fix the value of $g_{-}$ in the following. The choice to fix $g_{-}$ is further supported by the observation that, for finite $\Delta v>\Delta v^{*}$, we verified that keeping $0<g_{-}\lesssim 0.5$ does not affect the stripe orientation. Additionally, our simulations show that increasing $g_{0}$ consistently destabilizes perpendicular stripes, moving the system away from the phenomenology of the lattice models. Accordingly, we set $g_{-}=0.5$ and $g_{0}=0$ in what follows  [^13] Note that it is not strictly necessary to drastically set $g_{0}=0$; any small value $g_{0}\approx 0$ would yield similar results. However, since no particular value is preferred, we choose zero for simplicity and computational efficiency.. In other words, we focus the remainder of our studies on Regime 2 shown in Fig. 10. In this regime, we explore the effect of varying $\rho$ [which will vary the linear terms shown in Eq. (B)] and $\Delta v$ (or equivalently $\delta$), and find that increasing either parameter also increases order along the drive direction. Conversely, when $\rho$ and $\Delta v$ are close to the thresholds $(\rho_{c}^{\mathrm{FT}},\Delta v^{*})$, we find evidence of long-wavelength (Eckhaus-like) instabilities within the perpendicular stripe regime, which act as a mechanism to reduce the number of stripes, similarly to what occurs in the FLM. Based on this, we identify the large and near-threshold parameter regimes as the regular and irregular stripe regimes, respectively. The resemblance goes beyond visual inspection of the configurations: we also detect the clear signatures of the different regimes in the charge SF, with the irregular stripe regime [Fig. 11(a)] showing a broad peak near the origin, whereas the regular stripe regime [Fig. 11(b)] displays a sharp, Dirac delta-like peak accompanied by pronounced higher-order harmonics. However, unlike the FLM irregular stripe regime, we observe weak harmonics around the peak and secondary maxima at $q_{x}>q^{*}$ for $q_{y}\neq 0$, indicating stripes that are more ordered along $x$ yet weakly modulated along  [^14] In fact, this vertical modulation seems to be a feature of the equations that is already present if $g_{\pm,0}=0$; see, e.g., Fig. 10  (1). $y$. Additionally, in the regular stripe regime, the continuum model further accentuates ordering: the SF is heavily localized at $q_{y}=0$, with modes at any $q_{y}\neq 0$ appearing almost as background noise; the resulting spectrum is strongly peaked and displays deep dips, resembling Bragg peaks in a diffraction grating that are consistent with a highly ordered stripe pattern.
@@
For example, comparing with the results in Table 1, the ratio of peak values $S_{-}(q^{*},q_{y})$ for $q_{y}L/2\pi=0,1$ in the continuum model using $\delta=1$ is $S_{-}\left(q^{\ast},0\right)/S_{-}\left(q^{\ast},0.05\right)\simeq 52$ in the irregular stripe regime ($\rho=0.4$) and $S_{-}\left(q^{\ast},0\right)/S_{-}\left(q^{\ast},0.05\right)\simeq 2700$ in the regular stripe regime ($\rho=0.6$) (nearly an order of magnitude larger than the FLM!). Table 2 lists the peak values for the first three $q_{y}$ modes available in these two cases. In addition, we note that the density at which the regular stripe regime emerges in the continuum model appears to be shifted relative to the FLM: regular stripes are observed already at $\rho=0.6$ for $\delta=1$ (whereas in the FLM, $\rho_{u}\simeq 0.65$). However, note that $\rho$-dependence in the continuum model only refers to the GE results for the linear coefficients [see Eq. (B)]. These assumptions are violated at high densities and we generally expect the location of the transition to be shifted due to the nonlinearities and noise.
@@
Figure 11: Regimes of irregular stripes and regular stripes found by numerical integration of Eq. (III), analogous to the ordered regimes of the FLM illustrated in Fig. 5. We use $\delta=1$, maintaining $\lambda=1,\,g_{0}=0$ and $g_{-}=0.5$. (a) $\rho=0.4$. (b) $\rho=0.6$, with linear coefficients fixed by Eqs. (B). Charge fluctuation (left panels) and density fluctuation (center panels) of typical steady-state field configurations. Right panels: charge fluctuation structure factor, $S_{-}(q_{x},q_{y})$, as a function of $q_{x}$ for $q_{y}\,L/2\pi=0,...,10$, in the two regimes. A movie related to this figure is available in [46].
@@
Finally, by fixing $\rho$ and varying the drive-dependent parameters $\Delta v$, $g_{0}$, and $g_{-}$, we uncover a wide range of regimes in Eqs. (III), many of which have not been reported in the previous lattice-based studies. These regimes typically occur when $|g_{0}\phi_{+}|\gg\Delta v$. To (partially) understand the origin of these new regimes, we reorganize the drive-dependent terms proportional to $\partial_{x}\phi_{-}$ in Eq. (III) into the combination
@@
If $|g_{0}\phi_{+}|\gg\Delta v$, we see from Eq. (17) that the flow direction of the CF is strongly dependent on the sign of $\phi_{+}$, facilitating the formation of complex flow patterns.
Fig. 13 shows snapshots taken at long times $t=10^{6}$ that illustrate the variety of new behaviors: regular [Fig. 13(a)] and irregular [Fig. 13(b)] stripes aligned with the drive, spatio-temporal chaos [Fig. 13(c)] and phase separation into bubbles modulated by shock-like envelopes [Fig. 13(d)]. Because these regimes do not appear in either the FLM or the DWRLG, we have restricted our primary analysis to parameter regions that better reflect those models. Nevertheless, Fig. 13 clearly demonstrates that even a minimal set of continuum equations can generate a remarkably rich phenomenology. While a systematic exploration of how each term contributes to the various steady states would be valuable, such an endeavor is computationally demanding and falls beyond the scope of this study. It is worth noting that similar regimes have been observed in other contexts, such as parallel stripes in the KLS model [61] and spatio-temporal chaos in the active non-reciprocal Cahn-Hilliard model [25].
Figure 13: Charge fluctuation $\phi_{-}$ snapshots at $t=10^{6}$ of the different regimes found by numerical integration of Eq. (III). We use $\rho=0.7$ [fixing the coefficients in Eq. (B)], $\lambda=1$, and $\sigma=0.01$ in all panels. (a) Regular and (b) irregular stripes aligned with the drive. (c) Spatio-temporal chaos. (d) Phase separation into bubbles modulated by shock-like envelopes. Parameters: (a) $\Delta v=0.05$, $g_{0}=0.003$, $g_{-}=0.5$. (b) $\Delta v=0.211$, $g_{0}=0.3$, $g_{-}=0.5$. (c) $\Delta v=0.211$, $g_{0}=0.03$, $g_{-}=0.5$. (d) $\Delta v=0.05$, $g_{0}=0.1$, $g_{-}=0.5$. A movie related to this figure is available in [46].